ISOELECTRONIC HOMOLOGS AND ISOMERS: TROPOLONE, 5-AZATROPOLONE,1-H-AZEPINE-4,5-DIONE SADDLE-POINTS, AND IONS

\underline{RICHARD L. REDINGTON}, Department of Chemistry and Biochemistry, Texas Tech University, Lubbock Texas 79409.